# manvr3d: A Platform for Human-in-the-loop Cell Tracking in Virtual Reality

### v3

@@ 1 Introduction
The tracking process consists classically of two stages. In the first detection stage, cells are located in individual images or image stacks. In some cases, this might also include segmentation, in which cell shapes and boundaries are extracted in addition to cell positions. In the second linking step, individual cells are matched between successive images, allowing cell trajectories and lineage to be extracted. Tracking algorithms perform these tasks automatically on the input image. This scientific topic has received considerable attention and many tools are available today. Recent algorithms rely either on conventional image processing, statistical methods (like Gaussian mixture models [2]) or deep learning methods [16]. Evaluating their performance or training supervised deep learning models require ground truth annotations, which are extremely time-consuming to create due to the effort involved in manually annotating cells and their links across frames.
One solution to this problem is to use sparse annotations combined with incremental human-in-the-loop deep learning. The neural cell-tracking model learns from human feedback, provided in the form of iterative cycles of corrections to the model’s predictions. This is the approach taken by ELEPHANT [26]. However, with ELEPHANT, the human is constrained to a traditional mouse and keyboard interface and 2D display of 2D slices of the 3D data. This interface is not only slow but also error-prone, where the researcher might miss crucial spatial context.
@@ 2 Related work
Various software solutions exist for analyzing and visualizing biological data in 3D and/or VR environments. For cell lineage data, visualization and annotation solutions exist for both 2D [13, 22] and 3D [9, 23]. Neither of these support immersive rendering in VR. Leeuw et al. [3] address this gap by rendering cell trajectories in a CAVE [^1] Cave automatic virtual environment system with superimposed volume time-series. None of those solutions integrate the visualization aspects into a wider analysis platform and do not offer integration of a machine learning model.
Additionally, proprietary solutions like ConfocalVR [25] or syGlass [18] offer general quantification and measurement tools for biological image data, Arivis Pro VR222zeiss.com/microscopy/en/products/software/arivis-pro.html also offers features for cell tracking in volume data in VR. Arivis and syGlass have been utilized by Kaltenecker et al. [12] to annotate cell segmentations, which they then use to train their 3D deep learning model DELiVR for automated cell segmentation to speed up the annotation process. Note that in this work, no cell tracking is performed. Elor et al. developed the mixed reality environment BioLumin [4] to study the efficacy of crowd-sourced tissue annotation tasks, where the collected data is later used for training deep learning models. VR has been used in [5] to segment biological 3D time-series data, [28] visualized meshes resulting from the segmentation process, and [21] further annotate those meshes in VR. Again, no cell tracking is performed.
In contrast to the aforementioned solutions, we present an immersive open-source visualization platform for human-in-the-loop cell tracking with natural user interfaces to enable a streamlined and ergonomic process for both the creation of ground truth data and the final proofreading step. Our system is embedded into the Fiji [24] ecosystem, resulting in easier adoption and integration into existing pipelines.
@@ 3.1 Software ecosystem
manvr3d integrates with Fiji/ImageJ [24], a widely-used software package for both visualization and analysis of scientific and biological image data. For the purpose of this project, we rely on three existing Fiji plugins and frameworks:
  • Mastodon [20] as a platform for cell tracking, supporting the annotation of very large datasets. Mastodon uses BigDataViewer [19] as a backend to efficiently render large volumetric data as 2D slices. Tracking is performed manually using mouse and keyboard interaction, or semi-automatically with a difference-of-Gaussians filter for cell detection and either a Kalman filter or LAP linker [11] for cell linking.
  • ELEPHANT [26] extends Mastodon with an incremental deep learning model. The model is first trained on a sparse dataset, consisting only of a handful of manual annotations. In a proofreading loop, the user then corrects the predictions and iteratively trains the model again and again, quickly increasing the size of the training dataset. A second (optional) U-Net model is trained on optical flow prediction, using the existing cell links as training data, to guide the linking of spots between time points.
  • Sciview [6] and its underlying rendering framework scenery [8], to allow visualization of large volumetric data together with 3D meshes. Both tools support a variety of natural user interfaces, such as VR headsets, or eye tracking hardware.
By extending widely-used open-source software packages, we aim to broaden the appeal of our platform to users, and encourage researchers to extend upon it further. Within Mastodon, manvr3d acts as an extension and can be opened by selecting Window >>> New manvr3d view.
@@ 4.1 Cell Tracking with Handheld Controllers
Figure 3: Left: Scheme of the graph search algorithm that connects the closest local maxima in subsequent gaze rays following the first local maximum found in the first ray. Right: A collection of gaze rays across time, collected by following a moving cell with one’s eyes. We slide a simple Gauss kernel in the shape of [0.25,0.5,0.25]0.250.50.25[0.25,0.5,0.25][ 0.25 , 0.5 , 0.25 ] along each ray to smooth the signal and extract local maxima more effectively. Here we show the effect of various iterations of Gauss smoothing. It can be seen that no or low amounts of smoothing can lead to incorrectly extracted tracks, and starting from 4 iterations onward we are able to extract the correct cell track.
@@ 4.2 Cell Tracking via Eye Tracking
Tracking cells using gaze interaction has been explored in Bionic Tracking [7] by Günther et al. We incorporate this technique into our platform. While following a moving cell through the 3D volume data with their gaze, we record the user’s gaze directions and sample the volume at a uniform interval along each gaze ray. Finding the position of a cell along this gaze ray thus turns into a 1-dimensional problem, because we can assume that the first local maximum along the ray corresponds to the target cell. After calculating the local maxima for each ray, the track is reconstructed with a variant of the A* algorithm [27] that connects the closest local maxima from subsequent gaze rays, see Fig. 3.
To avoid the Midas touch problem [10], it is important to remove any visual distractions that could lead the user to unintentionally look away from their target cell. For this reason, we implemented a dual input approach for cell tracking by eye tracking: as soon as the user is ready to start tracking a cell with their eyes, they press the left trigger button. This starts playback of the dataset and the continuous collection of gaze directions and volume density samples along each ray. Once the user interrupts the tracking with the trigger button, or if the first time point is reached, the collected gaze rays and their sampled values are analyzed and subsequently sent to Mastodon. A collection of rays is plotted in Fig. 3, where each ray originates from the positive Y axis and extends downwards. Changes over time are plotted along the X axis.
@@ 5.2 Annotation performance
As stated before, performing a full user study is beyond the scope of this work. We instead performed two partial annotations (Table 2) of a Platynereis dataset and a Drosophila dataset [14], to compare annotation speeds between the Mastodon approach (2D), VR-based controller tracking (VR) and ELEPHANT (+DL). We measured the time taken to create 10 tracks each. We then used the pre-trained versatile ELEPHANT model and trained it over 5 epochs on the annotations created, and divided the final amount of tracks (104 for Platynereis and 74 for Drosophila) by the sum of training and prediction time. The training/prediction time is independent of the manual annotation method used.
@@ 6 Summary and future work
The ELEPHANT model provides uncertainty information to indicate the confidence of the network’s prediction. We plan to incorporate these data into the visualization process by coloring the spots and tracks accordingly, thus guiding the user towards areas of potentially higher error rate. Visualizing these uncertainty data is still an area of active research [1].
@@ A.1 Mastodon data structure
The data structure employed by Mastodon is highly efficient, with the entire directed graph stored internally as primitive byte arrays [20]. The reasons for this are vastly improved data access speed and memory efficiency compared to storing all vertices and edges of the graph as Java objects. One array stores all spots – the cell positions – and a second array stores all links that connect those spots. Both arrays reference each other using indices, and their content is accessed via proxy objects and iterators. A single spot stores data for the first incoming and outgoing link, respectively, as well as spot attributes like color and a covariance matrix that contains the spot’s scale and rotation. In addition to the indices for its source and target spots, each link also stores the indices for the next source or target link in the case of merge or split events.

### v4

@@ 1 Introduction
The tracking process consists classically of two stages. In the first detection stage, cells are located in individual images or image stacks. In some cases, this might also include segmentation, in which cell shapes and boundaries are extracted in addition to cell positions. In the second linking step, individual cells are matched between successive images, allowing cell trajectories and lineage to be extracted. Tracking algorithms perform these tasks automatically on the input image. This scientific topic has received considerable attention and many tools are available today. Recent algorithms rely either on conventional image processing, statistical methods (like Gaussian mixture models [amat_fast_2014]) or deep learning methods [deeplearningreview]. Evaluating their performance or training supervised deep learning models require ground truth annotations, which are extremely time-consuming to create due to the effort involved in manually annotating cells and their links across frames.
One solution to this problem is to use sparse annotations combined with incremental human-in-the-loop deep learning. The neural cell-tracking model learns from human feedback, provided in the form of iterative cycles of corrections to the model’s predictions. This is the approach taken by ELEPHANT [elephant]. However, with ELEPHANT, the human is constrained to a traditional mouse and keyboard interface and 2D display of 2D slices of the 3D data. This interface is not only slow but also error-prone, where the researcher might miss crucial spatial context.
@@ 2 Related work
Various software solutions exist for analyzing and visualizing biological data in 3D and/or VR environments. For cell lineage data, visualization and annotation solutions exist for both 2D [lange_aardvark_2025, pretorius_cell_2015] and 3D [hong_lineaged_2022, salvador-martinez_celavi_2021]. Neither of these support immersive rendering in VR. Leeuw et al. [de_leeuw_visualization_2000] address this gap by rendering cell trajectories in a CAVE [^1] Cave automatic virtual environment system with superimposed volume time-series. None of those solutions integrate the visualization aspects into a wider analysis platform and do not offer integration of a machine learning model.
Additionally, proprietary solutions like ConfocalVR [stefani_confocalvr_2018] or syGlass [pidhorskyi_syglass_2018] offer general quantification and measurement tools for biological image data, Arivis Pro VR222zeiss.com/microscopy/en/products/software/arivis-pro.html also offers features for cell tracking in volume data in VR. Arivis and syGlass have been utilized by Kaltenecker et al. [kaltenecker_virtual_2024] to annotate cell segmentations, which they then use to train their 3D deep learning model DELiVR for automated cell segmentation to speed up the annotation process. Note that in this work, no cell tracking is performed. Elor et al. developed the mixed reality environment BioLumin [elor_biolumin_2022] to study the efficacy of crowd-sourced tissue annotation tasks, where the collected data is later used for training deep learning models. VR has been used in [guerinot_new_2022] to segment biological 3D time-series data, [zhang_4d_2023] visualized meshes resulting from the segmentation process, and [platt_micellanngelo_2023] further annotate those meshes in VR. Again, no cell tracking is performed.
In contrast to the aforementioned solutions, we present an immersive open-source visualization platform for human-in-the-loop cell tracking with natural user interfaces to enable a streamlined and ergonomic process for both the creation of ground truth data and the final proofreading step. Our system is embedded into the Fiji [fiji] ecosystem, resulting in easier adoption and integration into existing pipelines.
@@ 3.1 Software ecosystem
manvr3d integrates with Fiji/ImageJ [fiji], a widely-used software package for both visualization and analysis of scientific and biological image data. For the purpose of this project, we rely on three existing Fiji plugins and frameworks:
  • Mastodon [mastodon] as a platform for cell tracking, supporting the annotation of very large datasets. Mastodon uses BigDataViewer [pietzsch_bigdataviewer_2015] as a backend to efficiently render large volumetric data as 2D slices. Tracking is performed manually using mouse and keyboard interaction, or semi-automatically with a difference-of-Gaussians filter for cell detection and either a Kalman filter or LAP linker [jaqaman_robust_2008] for cell linking.
  • ELEPHANT [elephant] extends Mastodon with an incremental deep learning model. The model is first trained on a sparse dataset, consisting only of a handful of manual annotations. In a proofreading loop, the user then corrects the predictions and iteratively trains the model again and again, quickly increasing the size of the training dataset. A second (optional) U-Net model is trained on optical flow prediction, using the existing cell links as training data, to guide the linking of spots between time points.
  • Sciview [sciview] and its underlying rendering framework scenery [scenery], to allow visualization of large volumetric data together with 3D meshes. Both tools support a variety of natural user interfaces, such as VR headsets, or eye tracking hardware.
By extending widely-used open-source software packages, we aim to broaden the appeal of our platform to users, and encourage researchers to extend upon it further. Within Mastodon, manvr3d acts as an extension and can be opened by selecting Window >> New manvr3d view.
@@ 4.1 Cell Tracking with Handheld Controllers
Figure 3: Left: Scheme of the graph search algorithm that connects the closest local maxima in subsequent gaze rays following the first local maximum found in the first ray. Right: A collection of gaze rays across time, collected by following a moving cell with one’s eyes. We slide a simple Gauss kernel in the shape of $[0.25,0.5,0.25]$ along each ray to smooth the signal and extract local maxima more effectively. Here we show the effect of various iterations of Gauss smoothing. It can be seen that no or low amounts of smoothing can lead to incorrectly extracted tracks, and starting from 4 iterations onward we are able to extract the correct cell track.
@@ 4.2 Cell Tracking via Eye Tracking
Tracking cells using gaze interaction has been explored in Bionic Tracking [gunther_bionic_2020] by Günther et al. We incorporate this technique into our platform. While following a moving cell through the 3D volume data with their gaze, we record the user’s gaze directions and sample the volume at a uniform interval along each gaze ray. Finding the position of a cell along this gaze ray thus turns into a 1-dimensional problem, because we can assume that the first local maximum along the ray corresponds to the target cell. After calculating the local maxima for each ray, the track is reconstructed with a variant of the A* algorithm [sun_dynamic_2009] that connects the closest local maxima from subsequent gaze rays, see Fig. 3.
To avoid the Midas touch problem [jacob_eye_1995], it is important to remove any visual distractions that could lead the user to unintentionally look away from their target cell. For this reason, we implemented a dual input approach for cell tracking by eye tracking: as soon as the user is ready to start tracking a cell with their eyes, they press the left trigger button. This starts playback of the dataset and the continuous collection of gaze directions and volume density samples along each ray. Once the user interrupts the tracking with the trigger button, or if the first time point is reached, the collected gaze rays and their sampled values are analyzed and subsequently sent to Mastodon. A collection of rays is plotted in Fig. 3, where each ray originates from the positive Y axis and extends downwards. Changes over time are plotted along the X axis.
@@ 5.2 Annotation performance
As stated before, performing a full user study is beyond the scope of this work. We instead performed two partial annotations (Table 2) of a Platynereis dataset and a Drosophila dataset [drosophila-dataset], to compare annotation speeds between the Mastodon approach (2D), VR-based controller tracking (VR) and ELEPHANT (+DL). We measured the time taken to create 10 tracks each. We then used the pre-trained versatile ELEPHANT model and trained it over 5 epochs on the annotations created, and divided the final amount of tracks (104 for Platynereis and 74 for Drosophila) by the sum of training and prediction time. The training/prediction time is independent of the manual annotation method used.
@@ 6 Summary and future work
The ELEPHANT model provides uncertainty information to indicate the confidence of the network’s prediction. We plan to incorporate these data into the visualization process by coloring the spots and tracks accordingly, thus guiding the user towards areas of potentially higher error rate. Visualizing these uncertainty data is still an area of active research [uncertainty].
@@ A.1 Mastodon data structure
The data structure employed by Mastodon is highly efficient, with the entire directed graph stored internally as primitive byte arrays [mastodon]. The reasons for this are vastly improved data access speed and memory efficiency compared to storing all vertices and edges of the graph as Java objects. One array stores all spots – the cell positions – and a second array stores all links that connect those spots. Both arrays reference each other using indices, and their content is accessed via proxy objects and iterators. A single spot stores data for the first incoming and outgoing link, respectively, as well as spot attributes like color and a covariance matrix that contains the spot’s scale and rotation. In addition to the indices for its source and target spots, each link also stores the indices for the next source or target link in the case of merge or split events.